A standard 1 kilogram weight is a cylinder 41.5 mm in height and 55.0 mm in diameter. What is the density of the material?
Answer #1

Mass of the cylinder = M = 1 kg

Density of the material = \rho

Diameter of the cylinder = D = 55 mm = 0.055 m

Radius of the cylinder = R = D/2 = 0.0275 m

Height of the cylinder = H = 41.5 mm = 0.0415 m

Volume of the cylinder = V

V = \piR2H

M = \rhoV

M = \rho\piR2H

1 = \rho\pi(0.0275)2(0.0415)

\rho = 10142.3 kg/m3

Density of the material = 10142.3 kg/m3